# Cellular, Cell-less, and Everything in Between: A Unified Framework for Utility Region Analysis in Wireless Networks

### v2

@@ Abstract
We introduce a unified framework for analyzing utility regions of wireless networks, with a focus on the signal-to-interference-noise-ratio (SINR) and achievable rate regions. The framework provides valuable insights into interference patterns of modern network architectures, such as cell-less and extremely large MIMO networks, and it generalizes existing characterizations of the weak Pareto boundary. A central contribution is the derivation of sufficient conditions that guarantee convexity of the utility regions. Convexity is an important property because it ensures that time sharing (or user grouping) cannot simultaneously increase the utility of all users when the network operates on the weak Pareto boundary. These sufficient conditions also have two key implications. First, they identify a family of (weighted) sum-rate maximization problems that are inherently convex without any variable transformations, thus paving the way for the development of efficient, provably optimal solvers for this family. Second, they provide a rigorous justification for formulating sum-rate maximization problems directly in terms of achievable rates, rather than SINR levels. Our theoretical insights also motivate an alternative to the concept of favorable propagation in the massive MIMO literature – one that explicitly accounts for self-interference and the beamforming strategy.
Index Terms:
Cell-less/massive MIMO, interference management, achievable SINR and rate regions
@@ I Introduction
In wireless networks with users competing for shared resources, system designers must balance fairness with overall network efficiency. To this end, the performance of each user is evaluated via a utility function, and this function guides resource allocation schemes to achieve the above balance. Two common design criteria for selecting efficient operating points are (weighted) max-min fairness [1, 2, 3, 4, 5][6, Ch. 3, Ch. 7][7, Ch.5.3.2][8, Theorem 7.1] and (weighted) sum utility maximization [9, 10, 11, 12], and these criteria often use the signal-to-interference-plus-noise ratio (SINR) or inner bounds of the Shannon capacity as the utility function.
One of the main advantages of the max-min fairness criterion is that it often leads to optimization problems that, although nonconvex in general, can be solved with efficient fixed point algorithms [1]. In addition, this criterion is particularly suitable for networks designed to ensure uniform service-of-quality across all users, as in cell-less or user-centric networks. However, user utilities can often be further improved through time sharing (link scheduling/user grouping) [13, Sect. 5.2.1] because, for mathematical tractability, the utilities are typically based on bounds of the capacity region, making the achievable utility region potentially nonconvex. While time sharing allows for greater flexibility in improving utilities, it also introduces a challenging combinatorial aspect to resource allocation. This aspect undermines one of the primary benefits of the max-min fairness criterion without time sharing: the availability of efficient algorithms able to compute global optima. In contrast, operating points that maximize the sum utility (typically the achievable rates) cannot be improved via time sharing. However, sum-rate maximization is NP-hard in general [14], so it is often addressed with fast heuristics without any guarantees of obtaining global optima [15].
As discussed in the next subsection, the above challenges are eliminated if the feasible utility region is convex. In this scenario, time sharing is unnecessary in max-min utility optimization problems, and fast iterative algorithms that provably converge to global optima can potentially be devised for sum-rate maximization problems. However, to date, existing methods for identifying convex utility regions or tractable sum-rate maximization problems rely on assumptions that are difficult to satisfy in practice. They include equal interference-plus-noise for all users in the network [12, Sect. V.A], symmetric interference patterns [13, Ch. 5.4.4], no interference, or low transmit power [11][9, Sect. IV.B]. Consequently, there is a need to identify weaker, more practical assumptions.
@@ I-A Time sharing in resource allocation
To describe the importance of time sharing, consider the uplink of a network with $N\in{\mathbb{N}}$ users represented by the set $\mathcal{N}=\{1,\ldots,N\}$, and let ${\bm{p}}=(p_{1},\ldots,p_{N})\in{\mathbb{R}}_{+}^{N}:=[0,~\infty[^{N}$ denote the transmit power vector, where $p_{n}$ represents the transmit power of user $n$. For each user $n\in\mathcal{N}$, we assign a utility function $u_{n}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ that maps a given power vector ${\bm{p}}\in{\mathbb{R}}_{+}^{N}$ to the quality-of-service experienced by the user.  [^1] These utilities take only the power vector as an argument, but other parameters to be optimized, such as the choice of beamformers and the user-base station assignment, can be implicitly considered in the definition of $u_{n}$. For concrete examples, see Sect. III and [16, 17, 18, 19, 20, 21]. In many wireless problems, which include cell-less systems, massive MIMO, and traditional cellular systems as particular instances, the utility functions, such as the SINR or the achievable rates, can be unified as follows [16, 20]:
@@
where $(t_{n}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++})_{n\in\mathcal{N}}$ are standard interference functions, in the sense defined in [22], and ${\mathbb{R}}_{++}:=]0,\infty[$. In real systems, the power vector is typically constrained to a set of the type $\mathcal{P}:=\{{\bm{p}}\in{\mathbb{R}}_{+}^{N}\mid\|{\bm{p}}\|\leq p_{\mathrm{max}}\}$, where $\|\cdot\|$ is a monotone norm (as defined in Sect. II-A) and $p_{\mathrm{max}}>0$ is the maximum transmit power. In this case, given a power vector ${\bm{p}}^{\star}\in\mathcal{P}$, we say that the network operates in an efficiency operating point, in the weak Pareto sense, if there is no other power vector that strictly increases the utility of every user in the system; i.e., the set $\{{\bm{p}}\in\mathcal{P}\mid(\forall n\in\mathcal{N})~u_{n}({\bm{p}})>u_{n}({\bm{p}}^{\star})\}$ is empty. With utilities having the structure in (1), efficient operating points can be easily parametrized in terms of the argument ${\bm{p}}\in\mathcal{P}$ of the utility functions. More precisely, the network with utilities $(u_{1}({\bm{p}}),\ldots,u_{N}({\bm{p}}))$ are efficient in the weak Pareto sense for a given ${\bm{p}}\in\mathcal{P}$ if and only if $\|{\bm{p}}\|=p_{\mathrm{max}}$ [20]; i.e., the set of boundary points of the utility region is given by $\{(u_{1}({\bm{p}}),\ldots,u_{N}({\bm{p}}))\in{\mathbb{R}}^{N}_{+}\mid\|{\bm{p}}\|=p_{\mathrm{\max}}\}$.
@@
Efficiency, in the sense defined above, is not enough from a system design perspective because the selected power vector also has to ensure fairness among users. In particular, operating points guaranteeing both efficiency and fairness can be obtained with the following optimization problem, which can be solved with simple fixed point algorithms such as those described in [1, 23] (Fact 6(ii) in the Supplemental Material):
$\begin{array}[]{rl}\text{maximize}&\min\limits_{n\in\mathcal{N}}\omega_{n}~u_{n}({\bm{p}})\\ \text{subject to}&{\bm{p}}\in\mathcal{P},\end{array}$ (4)
@@
where $(\omega_{1},\ldots,\omega_{N})\in{\mathbb{R}}_{+}^{N}$ are weights (priorities) assigned to users, and these weights establish the notion of fairness among users. For example, if $\omega_{n}=1$ for every $n\in\mathcal{N}$, we obtain the standard concept of max-min fairness, which has seen widespread application in cell-less networks in recent years [6, Ch. 3, Ch. 7][21].222 Another common optimization criterion is sum-rate maximization, which can also be characterized as a solution to a problem of the form in (4) for appropriately chosen weights. However, finding weights leading to a power vector maximizing the sum rate is itself a challenging problem if the achievable rate region is nonconvex.
@@
The solution to (4) ensures that no other power vector strictly increases the utility of all users. However, with time-invariant utilities, the average utility of all users may be further improved if we also allow groups of users to be scheduled via time sharing. Whether strictly increasing the average utility of all users is possible with time sharing depends on the shape of the set $\mathcal{S}_{\mathcal{P}}$ of achievable utilities, defined by $\mathcal{S}_{\mathcal{P}}:=\{(u_{1}({\bm{p}}),\ldots,u_{N}({\bm{p}}))\in{\mathbb{R}}^{N}_{+}\mid{\bm{p}}\in\mathcal{P}\}$. With time sharing, the users can achieve, on average, any point in the convex hull of $\mathcal{S}_{\mathcal{P}}$, so there may be opportunities to increase simultaneously the utility of every user if $\mathcal{S}_{\mathcal{P}}$ is nonconvex. More specifically, consider $L\in{\mathbb{N}}$ given power vectors ${\bm{p}}_{1},\ldots,{\bm{p}}_{L}$ from the set $\mathcal{P}$. Let $\alpha_{l}\geq 0$ denote the fraction of time allocated to the $l$th power vector, with $\sum_{l=1}^{L}\alpha_{l}=1$. The resulting time-average utility $\bar{u}_{n}$ of user $n\in\mathcal{N}$ is then $\bar{u}_{n}=\sum_{l=1}^{L}\alpha_{l}\,u_{n}({\bm{p}}_{l}).$ Therefore, if $\mathcal{S}_{\mathcal{P}}$ is a convex set, we verify that $(\bar{u}_{1},\ldots,\bar{u}_{N})\in\mathcal{S}_{\mathcal{P}}$ because $(\forall l\in\{1,\ldots,L\})~(u_{1}({\bm{p}}_{l}),\ldots,u_{N}({\bm{p}}_{l}))\in\mathcal{S}_{\mathcal{P}}$, so there exists a fixed power vector ${\bm{p}}\in\mathcal{P}$ such that $(u_{1}({\bm{p}}),\ldots,u_{N}({\bm{p}}))=(\bar{u}_{1},\ldots,\bar{u}_{N})$. In contrast, if $\mathcal{S}_{\mathcal{P}}$ is nonconvex, we may have $(\bar{u}_{1},\ldots,\bar{u}_{N})\notin\mathcal{S}_{\mathcal{P}}$; i.e., the time-average utility vector $(\bar{u}_{1},\ldots,\bar{u}_{N})$ may lie outside $\mathcal{S}_{\mathcal{P}}$, meaning that this utility vector cannot be achieved by any single power vector ${\bm{p}}\in\mathcal{P\textasciiacute}$.
@@
For example, set the standard $l_{\infty}$-norm, denoted by $\|\cdot\|_{\infty}$, as the monotone norm in the set $\mathcal{P}$ of power constraints, and consider the hypothetical scenario with two users depicted in Fig. 1, which shows a feasible utility (e.g., rate) region $\mathcal{S}_{\mathcal{P}}$. Point $C$ in the figure is the traditional max-min fair operating point by assuming that only adjustments to the power vector are possible. By allowing time sharing (link scheduling), we can alternate between the power vectors ${\bm{p}}_{1}=(p_{\mathrm{max}},~0)\in\mathcal{P}$ and ${\bm{p}}_{2}=(0,~p_{\mathrm{max}})\in\mathcal{P}$, using each vector half of the time. With this scheme, the users achieve, on average, the max-min utility corresponding to Point D in Fig. 1, where the average utility of user $n\in\{1,2\}$ is $\frac{1}{2}u_{n}({\bm{p}}_{1})+\frac{1}{2}u_{n}({\bm{p}}_{2})$. As a result, the average utility of both users strictly increases, even though we only use power vectors from the set $\mathcal{P}$. In contrast, if the utility region were convex in this example, any weighted max-min fair utility point achievable with time sharing would be also achievable by solving (4) for the given weights. Therefore, convexity of the utility region serves as a certificate that time sharing offers no additional benefit.
@@ I-B Main contributions
The first objective of the present study is to derive conditions that guarantee convexity of the SINR region, with and without power constraints. We also prove that our sufficient conditions for convexity of the SINR region imply convexity of the achievable rate region obtained by treating interference as noise. Our results cover a range of resource allocation problems in a unified way, including those described in [6, Ch. 3, Ch. 7] for cell-less networks, in [7, Ch. 5.3.2] [8, Theorem 7.1] for massive MIMO systems, and in [13, Ch. 5.6.3] for traditional wireless networks, among others.
As direct consequences of our results, we identify specific problem instances where sum-rate maximization considering modern interference models can be posed as a convex optimization problem without any changes in variables, and where the performance of network configurations obtained by solving max-min problems cannot be further enhanced through time sharing. We also provide a strong justification for using directly the rates as the optimization variables in sum-rate optimization problems, rather than SINR levels. In addition, we resolve a conjecture in [13, Sect. 5.4.4][24], derive an alternative parametrization of the weak Pareto boundary of the SINR and rate regions in terms of the spectral radius of nonlinear mappings, and introduce a new concept of channel compatibility among users. Our notion of compatibility overcomes limitations of the standard definition of favorable propagation in the massive MIMO literature.

### v3

@@ Abstract
We introduce a unified framework for analyzing utility regions of wireless networks, with a focus on signal-to-interference-plus-noise-ratio (SINR) and achievable rate regions. The framework provides valuable insights into interference patterns of modern network architectures, including extremely large MIMO and cell-less networks. A central contribution is a simple characterization of feasible utility regions using the concept of spectral radius of nonlinear mappings. This characterization provides a powerful mathematical tool for wireless system design and analysis. For example, it allows us to generalize existing characterizations of the weak Pareto boundary using compact notation. It also allows us to derive tractable sufficient conditions for the identification of convex utility regions. This property is particularly important because, on the weak Pareto boundary, it guarantees that time sharing (or user grouping) cannot simultaneously improve the utilities of all users. Beyond geometrical insights, these sufficient conditions have two key implications. First, they identify a family of (weighted) sum-rate maximization problems that are inherently convex, thus paving the way for the development of efficient, provably optimal solvers for this family. Second, they provide justification for formulating sum-rate maximization problems directly in terms of achievable rates, rather than SINR levels. Our theoretical insights also motivate an alternative to the concept of favorable propagation in the massive MIMO literature – one that explicitly accounts for self-interference and the beamforming strategy.
@@ I Introduction
In wireless networks with users competing for shared resources, system designers must balance fairness with overall network efficiency. To this end, the performance of each user is evaluated via a utility function, and this function guides resource allocation toward achieving this balance. Two common design criteria for selecting operating points are (weighted) max-min fairness [29, 9, 39, 26, 12][15, Ch. 3, Ch. 7][23, Ch.5.3.2][2, Theorem 7.1] and (weighted) sum utility maximization [37, 38, 44, 14], and these criteria often use the signal-to-interference-plus-noise ratio (SINR) or inner bounds of the Shannon capacity as the utility function.
One of the main advantages of the max-min fairness criterion is that it often leads to optimization problems that, although nonconvex in general, can be solved with efficient fixed point algorithms [29]. In addition, this criterion is particularly suitable for networks designed to ensure uniform quality of service across all users, as in cell-less or user-centric networks. However, user utilities can often be further improved through time sharing (link scheduling/user grouping) [35, Sect. 5.2.1] because, for mathematical tractability, the utilities are typically based on capacity-region bounds or surrogate expressions based on the SINR, making the achievable utility region potentially nonconvex. While time sharing allows for greater flexibility in improving utilities, it also introduces a challenging combinatorial aspect to resource allocation. This aspect undermines one of the primary benefits of the max-min fairness criterion without time sharing: the availability of efficient algorithms able to compute global optima. In contrast, operating points that maximize the sum utility (typically the achievable rates) cannot be improved via time sharing. However, sum-rate maximization is NP-hard in general [22], so it is often addressed with fast heuristics without any guarantees of obtaining global optima [41, 25].
As discussed in the next subsection, the above challenges are eliminated if the feasible utility region is convex. In this scenario, time sharing is unnecessary in max-min utility optimization problems, and fast iterative algorithms that provably converge to global optima can potentially be devised for sum-rate maximization problems. However, to date, existing methods for identifying convex utility regions or tractable sum-rate maximization problems rely on assumptions that are difficult to satisfy in practice. They include equal interference-plus-noise for all users in the network [14, Sect. V.A], symmetric interference patterns [35, Ch. 5.4.4], no interference, or low transmit power [44][37, Sect. IV.B]. Consequently, there is a need to identify weaker, more practical assumptions.
@@ I-A Time sharing in resource allocation
To illustrate how the geometry of the utility region governs the need for time sharing in resource allocation, consider the uplink of a network with $N\in{\mathbb{N}}$ users represented by the set $\mathcal{N}=\{1,\ldots,N\}$, and let ${\bm{p}}=(p_{1},\ldots,p_{N})\in{\mathbb{R}}_{+}^{N}:=[0,~\infty[^{N}$ denote the transmit power vector, where $p_{n}$ represents the transmit power of user $n$. For each user $n\in\mathcal{N}$, we assign a utility function $U_{n}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ that maps a given power vector ${\bm{p}}\in{\mathbb{R}}_{+}^{N}$ to the quality-of-service experienced by the user.  [^1] These utilities take only the power vector as an argument, but other parameters to be optimized, such as the choice of beamformers and the user-base station assignment, can be implicitly considered in the definition of $U_{n}$. For concrete examples, see Sect. III and [5, 4, 33, 31, 10, 27]. In many wireless problems, which include cell-less systems, massive MIMO, and traditional cellular systems as particular instances, the utility functions, such as the SINR or the achievable rates, can be unified as follows [5, 10]:
@@
where $(t_{n}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++})_{n\in\mathcal{N}}$ are standard interference functions, in the sense defined in [42], and ${\mathbb{R}}_{++}:=]0,\infty[$. In real systems, the power vector is typically constrained to a set of the type $\mathcal{P}:=\{{\bm{p}}\in{\mathbb{R}}_{+}^{N}\mid\|{\bm{p}}\|\leq p_{\mathrm{max}}\}$, where $\|\cdot\|$ is a monotone norm (as defined in Sect. II-A) and $p_{\mathrm{max}}>0$ is the maximum transmit power. In this case, given a power vector ${\bm{p}}^{\star}\in\mathcal{P}$, we say that the network operates at an efficient operating point, in the weak Pareto sense, if there is no other power vector in $\mathcal{P}$ that strictly increases the utility of every user in the system; i.e., the set $\{{\bm{p}}\in\mathcal{P}\mid(\forall n\in\mathcal{N})~U_{n}({\bm{p}})>U_{n}({\bm{p}}^{\star})\}$ is empty. With utilities having the structure in (1), efficient operating points can be easily parametrized in terms of the argument ${\bm{p}}\in\mathcal{P}$ of the utility functions. More precisely, under the constraint ${\bm{p}}\in\mathcal{P}$, the network with utilities $(U_{1}({\bm{p}}),\ldots,U_{N}({\bm{p}}))$ operates at an efficient point, in the weak Pareto sense, if and only if $\|{\bm{p}}\|=p_{\mathrm{max}}$ [10]; i.e., the weak Pareto boundary of the utility region is given by $\{(U_{1}({\bm{p}}),\ldots,U_{N}({\bm{p}}))\in{\mathbb{R}}^{N}_{+}\mid\|{\bm{p}}\|=p_{\mathrm{\max}}\text{ and }{\bm{p}}\in{\mathbb{R}}^{N}_{+}\}$.
@@
Efficiency, in the sense defined above, is not enough from a system design perspective because the selected power vector also has to ensure fairness among users. In particular, operating points guaranteeing both efficiency and fairness can be obtained with the following optimization problem, which can be solved with simple fixed point algorithms such as those described in Fact 6(ii) in the Supplemental Material [29, 9]:
$\begin{array}[]{rl}\mathrm{maximize}_{{\bm{p}}\in\mathcal{P}}&\min\limits_{n\in\mathcal{N}}\omega_{n}~U_{n}({\bm{p}}),\end{array}$ (3)
@@
where $(\omega_{1},\ldots,\omega_{N})\in{\mathbb{R}}_{++}^{N}$ are weights (priorities) assigned to users, and these weights establish the notion of fairness among users. For example, if $\omega_{n}=1$ for every $n\in\mathcal{N}$, we obtain the standard concept of max-min fairness, which has seen widespread application in cell-less networks in recent years [15, Ch. 3, Ch. 7][27]. [^2] Another common optimization criterion is sum-rate maximization, which can also be characterized as a solution to a problem of the form in (3) for appropriately chosen weights. However, finding weights leading to a power vector maximizing the sum rate is itself a challenging problem if the achievable rate region is nonconvex.
@@
The solution to (3) ensures that no other fixed power allocation strategy strictly increases the utility of all users. However, with time-invariant utilities, the average utility of every user may be improved by allowing time sharing; i.e., scheduling different groups of users (or operating points) over time. Crucially, whether time sharing can increase the utility of every user relative to the fixed power allocation obtained from (3) – regardless of the particular fairness weights adopted in (3) – is governed by the geometry of the achievable utility region $\mathcal{S}_{\mathcal{P}}:=\{(U_{1}({\bm{p}}),\ldots,U_{N}({\bm{p}}))\in{\mathbb{R}}^{N}_{+}\mid{\bm{p}}\in\mathcal{P}\}$. To illustrate this link, consider $L\in{\mathbb{N}}$ given power vectors ${\bm{p}}_{1},\ldots,{\bm{p}}_{L}$ from the set $\mathcal{P}$. Let $\alpha_{l}\geq 0$ denote the fraction of time allocated to the $l$th power vector, with $\sum_{l=1}^{L}\alpha_{l}=1$. The resulting time-average utility $\bar{U}_{n}$ of user $n\in\mathcal{N}$ is then $\bar{U}_{n}=\sum_{l=1}^{L}\alpha_{l}\,U_{n}({\bm{p}}_{l}).$ Therefore, the time-average utility vector $(\bar{U}_{1},\ldots,\bar{U}_{N})$ is a convex combination of points in $\mathcal{S}_{\mathcal{P}}$. If $\mathcal{S}_{\mathcal{P}}$ is convex, then $(\bar{U}_{1},\ldots,\bar{U}_{N})\in\mathcal{S}_{\mathcal{P}}$, so there exists a single power vector ${\bm{p}}\in\mathcal{P}$ such that $(U_{1}({\bm{p}}),\ldots,U_{N}({\bm{p}}))=(\bar{U}_{1},\ldots,\bar{U}_{N})$; i.e., time sharing does not enlarge the achievable utility region. In contrast, if $\mathcal{S}_{\mathcal{P}}$ is nonconvex, the time-average utility vector $(\bar{U}_{1},\ldots,\bar{U}_{N})$ may lie outside $\mathcal{S}_{\mathcal{P}}$, meaning that this utility vector cannot be achieved by any single power vector ${\bm{p}}\in\mathcal{P}$.
@@
For concreteness, set the standard $l_{\infty}$-norm, denoted by $\|\cdot\|_{\infty}$, as the monotone norm in the set $\mathcal{P}$ of power constraints, and consider the hypothetical scenario with two users depicted in Fig. 1, which shows a feasible utility (e.g., rate) region $\mathcal{S}_{\mathcal{P}}$. Point $C$ in the figure is the traditional max-min fair operating point by assuming that only adjustments to the power vector are possible. By allowing time sharing (link scheduling), we can alternate between the power vectors ${\bm{p}}_{1}=(p_{\mathrm{max}},~0)\in\mathcal{P}$ and ${\bm{p}}_{2}=(0,~p_{\mathrm{max}})\in\mathcal{P}$, using each vector half of the time. With this scheme, the users achieve, on average, the max-min utility corresponding to Point D in Fig. 1, where the average utility of user $n\in\{1,2\}$ is $\frac{1}{2}U_{n}({\bm{p}}_{1})+\frac{1}{2}U_{n}({\bm{p}}_{2})$. As a result, the average utility of both users strictly increases, even though we only use power vectors from the set $\mathcal{P}$. In contrast, if the utility region were convex in this example, then, for the given weights defining the fairness criterion, no weighted max-min fair utility point obtainable via time sharing would strictly improve upon the utilities achieved by the solution to (3). Therefore, convexity of the utility region serves as a certificate that time sharing offers no additional benefit.
@@ I-B Main contributions
The first objective of this study is to link solutions to max-min problems to nonlinear eigenpairs of general interference mappings (see Sect. II-A for the definitions), thereby providing a concise description of the solutions (Remark 1). The key step is a new result establishing a correspondence between conditional eigenpairs of standard interference mappings with eigenpairs of related general interference mappings (Proposition 1). Building on this correspondence, we derive compact representations of the feasible SINR and rate regions of wireless networks, both with and without power constraints, in terms of the concept of nonlinear spectral radius – see (27), (29), (31), and (32). A central implication of this characterization is that it allows us to derive sufficient conditions – based on inverse Z-matrices – that guarantee convexity of the SINR region, and, consequently, convexity of the rate region (Corollary 1 and Corollary 2). As discussed above, convexity has two important practical consequences. First, when these regions are convex, the performance attained by solutions to max-min problems cannot be improved via time sharing. Second, as shown in Sect. IV-E, sum-rate maximization can be formulated as a tractable convex optimization problem. Our framework covers a broad class of resource-allocation models in a unified manner, including those in [15, Ch. 3, Ch. 7] for cell-less networks, [23, Ch. 5.3.2] and [2, Theorem 7.1] for massive MIMO systems, and [35, Ch. 5.6.3] for traditional wireless networks, among others.
As further consequences of our results, we obtain a new characterization of the weak Pareto boundary of the SINR and rate regions in terms of the spectral radius of nonlinear mappings (Proposition 5), resolve the conjecture in [35, Sect. 5.4.4][36] (Sect. IV-D), and introduce a new concept of channel compatibility among users that overcomes some limitations of the standard definition of favorable propagation in the massive MIMO literature (Sect. IV-B).